Find the coordinates of B. M is the midpoint of line AB.
A(7,5), M(9,7)

Answer by stanbon(75887)

Find the coordinates of B. M is the midpoint of line AB.
A(7,5), M(9,7)
-------------------------
Let B have coordinates (x,y)
---------------
If M is the midpoint (x+7)/2 = 9 ; x+7 = 18 ; x = 11
----------------
If M is the midpoint (y+5)/2 = 7 ; y+5 = 14 ; y = 9
-------------------------
B has coordinates (11,9)
===========================
Cheers,
Stan H.